OpenStudy (anonymous):

y=x^2+2 y=-x+4 SOMEONE PLEASE HELP ME ON THIS

OpenStudy (anonymous):

Equate the two equations, and then move everythign to one side so you have x^2+x-2. That factors into (x-1)^2=0, therefore x=1

OpenStudy (anonymous):

Y=y, so x^2+2=-x+4 x^2+x-2=0 (x-1)(x+2)=0 x=1 or -2 Y=-x+4 y=-1+4. Or y=-(-2)+4 y=3 or y=6 so, x=1, y=3 or x=-2, y=6

OpenStudy (anonymous):

@spaceknight ... after you get that 1 , do i plug that into x to get y ?

OpenStudy (anonymous):

&lotsobooks omg thanks . can you like tell me the steps on how you do it

OpenStudy (anonymous):

lol oops i factored wrong sorry just listen to lotsobooks

OpenStudy (anonymous):

i already solved this problem in the one before...

OpenStudy (anonymous):

ok so can i ask you guys another question ? i need help on my entire algebra2 homework

OpenStudy (anonymous):

you should try doing your own homework

OpenStudy (anonymous):

&spaceknight i have been trying . if i knew how to do thiss i wouldntt be on my laptop asking for help . my teacher is the kind that doesnt like to explain too much .... I NEED HELP

OpenStudy (anonymous):

Set both equations equal to one another. Move all of the terms to the left side of the equation by adding or subtracting the terms on both sides of the equation. Combine like terms. Since it is a quadratic equation, you can factor it - use FOIL backwards, sort of. Once you have the 2 possible values of x, then, you can substitute them back inti either of the original equations - I picked -x+4 since it is the easiest. Then, calculate the values of y. Remember to keep the x,y pairs together since they only go with one another.

OpenStudy (anonymous):

see i never got instructions like that. ok how about this . im about to show you my next problem. im going to try and solve it by myself to see if i can do it but can you still help me

OpenStudy (anonymous):

?

OpenStudy (anonymous):

y=6x^2+5x+3 y==x+3

OpenStudy (anonymous):

\[6x ^{2}+5x+3=-x+3\]

OpenStudy (anonymous):

So far, so good.

OpenStudy (anonymous):

ok so it will go like this i thinkk 6x^2+5x+3=-x+3 +x-3 +x-3 hold on i think i messed up

OpenStudy (anonymous):

i got 6x^2+6x=0 when i subtracted

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

so what do i do now ?

OpenStudy (anonymous):

In this case, it is easiest to subtract 6x from both sides.

OpenStudy (anonymous):

then \[6x ^{2} = -6x ?\]

OpenStudy (anonymous):

Right. Then, divide both sides by 6.

OpenStudy (anonymous):

ok so then that is x^2 = -1x or x^2 = -x ?

OpenStudy (anonymous):

Right. Now, divide both sides by x.

OpenStudy (anonymous):

ok so then after that x = -1 ?

OpenStudy (anonymous):

Right. Plug -1 into one of the original equations and solve for y.

OpenStudy (anonymous):

so whenever i plug it into the second equation how do i do it ? do i ignore the negative there already is ?

OpenStudy (anonymous):

No, keep the negative. -(-1)+3=

OpenStudy (anonymous):

wouldnt that 1 just turn into a positive ?

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

ok so it would be 1+3 so y = 4 ?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

We missed a second solution when I had you subtract 6x from both sides.

OpenStudy (anonymous):

you are AWESOMEEEE !!!! :p thanks for helping meee

OpenStudy (anonymous):

what step did we miss ?

OpenStudy (anonymous):

Not a step so much as another solution. At 6x^2+6x=0, it would have been better to factor 6x out. So, 6x (x+1)=0 x=-1 or x=0. then, y=4 or y=3.

OpenStudy (anonymous):

You're welcome! Have fun!

OpenStudy (anonymous):

:)